### v1

# Simulated Laser Cooling and Magneto-Optical Trapping of Group IV Atoms

Abstract
We present a scheme for laser cooling and magneto-optical trapping of the Group IV (a.k.a. Group 14 or tetrel) atoms silicon (Si), germanium (Ge), tin (Sn), and lead (Pb). These elements each possess a strong Type-II transition ($J\rightarrow J^{\prime}=J-1$) between the metastable $s^{2}p^{2}\,{}^{3}P_{1}$ state and the excited $s^{2}ps^{\prime}\,{}^{3}P_{0}^{o}$ state at an accessible laser wavelength, making them amenable to laser cooling and trapping. We focus on the application of this scheme to Sn, which has several features that make it attractive for precision measurement applications. We perform numerical simulations of atomic beam slowing, capture into a magneto-optical trap (MOT), and subsequent sub-Doppler cooling and compression in a blue-detuned MOT of Sn atoms. We also discuss a realistic experimental setup for realizing a high phase-space density sample of Sn atoms.
I Introduction
[…]
Laser cooling and magneto-optical trapping have also been applied to molecules, unlocking a new class of quantum matter to explore. Molecular structure dictates that optical cycling must operate on a Type-II transition ($F^{\prime}=F-1$ or $F^{\prime}=F$) for rotational closure. To date, several molecular species have been laser cooled and trapped in a MOT [28, 29, 30, 31, 32, 33, 34]. Among these molecules, CaF, SrF, YO, CaOH, and SrOH have been further cooled to ultracold temperatures ($\sim\!5-100\,\mu\text{K}$) and loaded into conservative optical dipole traps (ODTs) [35, 36, 37, 38, 39]. These developments indicate that Type-II transitions are well-suited for use in various quantum science applications.
Adding a new class of elements to the list of laser-coolable species would constitute another leap in the quest to control and manipulate quantum matter. In this work, we present numerical simulations to verify a scheme for laser cooling and trapping of Group IV elements — silicon (Si), germanium (Ge), tin (Sn), and lead (Pb) — using a Type-II transition. Among these elements, previous attempts have been made to study pathways towards producing ultracold gases of Si atoms [40, 41, 42, 43], but a Si MOT has not been realized to date. Our work should provide a general path forward for laser cooling and trapping of Group IV elements for applications in quantum science and precision measurement.
II Laser Cooling and Trapping Scheme
II.1 Structure of Group IV Atoms
In their ground state, all Group IV elements have four valence electrons in a $s^{2}p^{2}$ atomic orbital configuration. Generically, there are five even parity states in the ground-state manifold, with terms $^{3}P_{0,1,2},{{}^{1}D}_{2}$, and $^{1}S_{0}$ in order of energy from lowest to highest. The lowest-lying odd parity excited state has configuration $s^{2}ps^{\prime}$ and term $^{3}P_{0}^{o}$. [^1] C and Si also have a $sp^{3}$ $^{5}S_{2}^{o}$ state below the $s^{2}ps^{\prime}$ $^{3}P^{o}$ states and above the $s^{2}p^{2}$ $^{1}S_{0}$ state, but this does not affect our discussion. All elements in this group have multiple spin-zero ($I=0$) even isotopes and several spinful odd isotopes; in this paper we focus only on the even isotopes. A level diagram of this structure is shown in Fig. 1.
[…]
We propose using the metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ to excited $s^{2}ps^{\prime}\,{{}^{3}}P_{0}^{o}$ transition in the Group IV elements as a pure Type-II optical cycling transition that is amenable to laser cooling and trapping. This transition is dipole-allowed with large natural linewidth. Moreover, it is completely closed by selection rules, meaning that no repumper laser is necessary. [^2] In Pb, the metastable state lifetime is short enough that a $^{3}P_{0}\rightarrow{{}^{3}}P_{1}$ repumper laser ($\lambda=1279$ nm) may be useful. Table 1 lists the relevant properties of this transition across all Group IV elements, as well as the metastable state lifetime. In the case of Si, Ge, Sn, and Pb, the transition wavelength can be produced with current ultraviolet (UV) laser technology. In the case of C, the transition wavelength is too short to currently be produced for laser cooling and trapping, so we do not consider it further in this work.
Table 1: Optical transition wavelength $\lambda_{c}$, natural linewidth $\Gamma$, and saturation intensity $I_{\text{sat}}$ of the $s^{2}p^{2}\,{{}^{3}}P_{1}\rightarrow s^{2}ps^{\prime}\,{{}^{3}}P_{0}^{o}$ transition proposed for use in laser cooling and trapping of Group IV elements [44]. The metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ state lifetime $\tau_{g}$ is also listed [44]. More details on calculating $\tau_{g}$ are given in the Supplemental Material (SM) [45].
Element $\lambda_{c}$ (nm) $\Gamma$ (MHz) $I_{\text{sat}}$ (mW/cm2) $\tau_{g}$ (s)
C 166 $2\pi\times 55$ 1572 $1.3\times 10^{7}$
Si 252 $2\pi\times 35$ 286 $1.2\times 10^{5}$
Ge 271 $2\pi\times 45$ 296 $3.2\times 10^{2}$
Sn 303 $2\pi\times 32$ 149 11
Pb 368 $2\pi\times 22$ 58 0.14
We note that a more conventional-seeming optical cycling transition is also present: $s^{2}p^{2}\,{}^{3}P_{2}$ to $sp^{3}\,{{}^{3}}D_{3}^{o}$ (for Si) or $s^{2}pd\,{{}^{3}}D_{3}^{o}$ (for Ge and Sn), i.e. the excited state is the lowest-lying odd-parity $D$ state. This was used to attempt laser cooling of Si ($\lambda=222\,\text{nm}$) [42]. This cycling transition is leaky, with an estimated branching ratio $^{3}D_{3}^{o}\rightarrow{{}^{1}}D_{2}$ of $\sim\!0.02\%-50\%$ depending on the element [44], necessitating a repumper laser to close the cycling transition. However, inclusion of this repumper creates a $\Lambda$-system, effectively making a Type-II cycling transition with two UV lasers on two different transitions. The cycling transition we propose is simpler and should lead to more effective laser cooling and trapping of Group IV elements.
In addition to supporting a Type-II transition, all Group IV elements possess four linearly independent clock transitions between the long-lived $s^{2}p^{2}$ states. Details on these transitions, which can only be of magnetic dipole (M1) or electric quadrupole (E2) type, are listed in Table 2.
Table 2: List of narrow linewidth clock transitions between $s^{2}p^{2}$ states in Group IV elements. Asterisk (*) denotes nominally spin-forbidden transitions. For the specific case of Sn, we list the wavelength $\lambda$ [44], calculated partial linewidth $\Gamma_{\text{p}}$ [46], and total linewidth $\Gamma_{\text{tot}}$.
Atomic Sn
State Transition Type $\lambda$ (nm) $\Gamma_{\rm p}$ (Hz) $\Gamma_{\rm tot}$ (Hz)
$^{3}P_{1}$ $\leftrightarrow\,^{3}P_{0}$ M1 5911 $2\pi\times 0.01$ $2\pi\times 0.01$
$^{3}P_{2}$ $\leftrightarrow\,^{3}P_{0}$ E2 2917 $2\pi\times 10^{-4}$ $2\pi\times 0.01$
$\leftrightarrow\,^{3}P_{1}$ M1/E2 5761 $2\pi\times 0.01$
$^{1}D_{2}$ $\leftrightarrow\,^{3}P_{0}$   E2* 1161 $2\pi\times 10^{-5}$ $2\pi\times 0.16$
$\leftrightarrow\,^{3}P_{1}$   M1/E2* 1445 $2\pi\times 0.07$
$\leftrightarrow\,^{3}P_{2}$   M1/E2* 1929 $2\pi\times 0.08$
$^{1}S_{0}$ $\leftrightarrow\,^{3}P_{1}$   M1* 646 $2\pi\times 1.13$ $2\pi\times 1.42$
$\leftrightarrow\,^{3}P_{2}$   E2* 728 $2\pi\times 0.11$
$\leftrightarrow\,^{1}D_{2}$ E2 1170 $2\pi\times 0.18$
II.2 Applications of Laser-Cooled Group IV Atoms
Ultracold atomic gases of Group IV elements have potential applications ranging from improved quantum sensing and novel quantum computing platforms to precision tests of fundamental physics and searches for physics beyond the Standard Model (BSM). Initial efforts to produce an ultracold atomic gas of Si atoms focused on envisioned applications to solid-state-based nuclear spin quantum computing [47, 48] and quantum sensing [49]. Resonant photoionization of ultracold radioactive $^{31}\text{Si}$ atoms in a MOT, followed by ion-guiding optics, would enable nm-scale precision implantation of $^{31}\text{Si}^{+}$ ions in a Si substrate [42, 50]. Subsequent $\beta$-decay of these ions ($\tau\sim\!3$ hrs) would create an array of $^{31}\text{P}^{+}$ ions doped in the Si substrate, realizing the building blocks of the Kane quantum computing architecture [47].
The plethora of accessible clock transitions in the ground $s^{2}p^{2}$ configuration can be utilized for a variety of precision measurement experiments. Sn is particularly attractive for such applications because it possesses a chain of seven stable, $I=0$ isotopes (the longest found in nature). The combination of many clock transitions and many $I=0$ isotopes makes laser-cooled and trapped Sn atoms a promising platform for probing non-linearity in King plot isotope shift measurements [51, 52]. Detection of non-linearity could indicate BSM physics manifesting as a new, light boson that interacts with electrons and nucleons [53]. Other charge states of Sn, such as $\text{Sn}^{2+}$ and $\text{Sn}^{7+}\text{-}\text{Sn}^{10+}$ ions, have been recently explored and also possess suitable clock transitions for such measurements [54, 55, 56, 57].
The properties of Group IV elements also make them attractive for the measurement of atomic parity violation (APV). Measurements of ratios of APV amplitudes in a long, stable isotope chain make it possible to cancel uncertainty from electronic structure factors that are difficult to compute from ab initio theory [58, 59, 60]. Clock transitions greatly improve energy resolution, and hence can provide a platform for more precise measurements of APV than in prior experiments [61]. Together, these features open a path to more sensitive probes of the nuclear weak charge [62, 63]. Alternatively, assumption of Standard Model values for weak interaction parameters enables measurements of APV along a chain of $I=0$ isotopes to precisely probe the “neutron skin” effect  [64, 65, 66]. Combined with other nuclear physics experiments, such measurements may provide critical insight into the equation of state of neutron-rich matter and establish a direct link between nuclei and neutron stars [67]. Isotopes with nonzero nuclear spin can also probe nuclear spin-dependent parity violation (NSD-PV), including contributions from the nuclear anapole moment where there is tension between nuclear theory and experimental measurements [68].
Finally, laser cooling of Group IV atoms introduces the opportunity to assemble ultracold, diatomic molecules built from Group IV and coinage metal atoms (Cu, Ag, Au), where laser cooling has been demonstrated (for Ag)  [22, 69]. This unique class of molecule was recently proposed as a promising candidate for next-generation precision measurements of the electron electric dipole moment (eEDM) [70].
II.3 Scheme for Laser Cooling and Trapping with the $^{3}P_{1}\rightarrow{{}^{3}P}_{0}^{o}$ Transition
[…]
The presence of dark Zeeman sublevels in the ground state necessitates a mechanism to destabilize such states during optical cycling, lest population accumulate there and cause cooling and trapping to cease [71]. It also leads to reduced photon scattering rate and optical force compared to conventional Type-I transitions. Zeeman slowing on a Type-II transition, though possible in principle, is difficult to realize [72, 73, 74]. Fortunately, these challenges have been successfully surmounted experimentally. In 1-D (e.g. for laser slowing), dark ground states are destabilized by rapidly switching the polarization of the slowing laser [32] or by applying a static magnetic field transverse to the slowing laser polarization [75]. In 3-D (e.g. in a MOT), destabilization is achieved by rapidly switching the laser polarizations [76] or by applying two laser frequencies with orthogonal polarizations [77, 78, 79]. The destablization rate must be on the order of the scattering rate for effective application of optical forces. For numerical simulations of laser slowing in this work, we focus only on the application of a transverse static magnetic field for remixing, while for simulations of the MOT, we focus only on the use of two laser frequencies with orthogonal polarizations.
We propose using an ablation-loaded cryogenic buffer gas beam (CBGB) — as is typically used in molecular laser cooling experiments [28, 29, 30, 80] — as a cold, bright source of atoms for a Group IV atomic laser cooling and trapping experiment. Typical experiments using CBGB sources report low initial velocities of $\sim\!100-200\,\text{m/s}$ [80]. Since laser slowing using a Type-II transition is inefficient compared to Type-I Zeeman slowers [75], the low initial forward velocity of atoms in CBGBs is attractive. Moreover, ablation loading leads to high initial temperatures ($T\sim\!4000\,\text{K}$) of the ablated species [81], which should naturally populate the metastable $^{3}P_{1}$ state that serves as the ground state of the optical cycling scheme. For example, we estimate that $\sim\!30\%$ of the population of Sn atoms will be in the $^{3}P_{1}$ state. Existing data and theory calculations [82, 83] suggest that the collisional quenching rate of metastable Group IV atoms in the buffer gas to the ground state is much slower than a typical extraction rate from the source cell. [^3] If collisional quenching does turn out to be a problem, it would be possible to optically pump population into the $^{3}P_{1}$ state.
In this work, we assume the source to have the same properties as the CBGB used in SrF laser cooling and trapping [84]. The buffer gas is helium at $T\sim\!4$ K, with flow rate $\mathcal{F}\sim\!5$ sccm. The initial longitudinal velocity distribution is roughly Gaussian, with mean longitudinal velocity $\bar{v}_{L}=140\,\text{m/s}$ and full width at half maximum (FWHM) $\Delta{v_{L}}=75\,\text{m/s}$. The initial transverse velocity distribution is also roughly Gaussian, with mean $\bar{v}_{T}=0\,\text{m/s}$ and FWHM $\Delta{v_{T}}=75\,\text{m/s}$. We assume a beam brightness across all isotopes of $\mathcal{B}\sim\!1.5\times 10^{12}$ atoms in the $^{3}P_{1}$ metastable state/steradian/pulse, commensurate with the measured brightness of Sr atoms in the SrF beam [85] after accounting for the metastable state population. We remark that using a thermal beam source of atoms may also be viable, but we do not explore this path here.
III Numerical Simulations of Optical Forces
In previous work from our group [74], simulation software based on solving the optical Bloch equations (OBEs) was developed to investigate molecular laser cooling and trapping. In this work, we adapt this software for the case of Group IV atomic laser cooling and trapping. Details of the formalism can be found in Refs. [74, 86]. In short, these simulations yield the acceleration of atoms due to optical forces. [^4] The acceleration is averaged over optical scattering cycles and randomized laser phases. Our simulations neglect particle-particle interactions and collisions. [^5] This precludes direct estimation of MOT loss rate, atom number, density, or optical thickness. MOTs realized using Type-II transitions are typically sparse enough that simulations under these assumptions still accurately capture the relevant physics of such MOTs.
To determine the force on atoms from external electromagnetic fields, we evolve the quantum state of the system until observables reach a periodic steady state [87]. We calculate the force using the Heisenberg equation of motion, averaged over one period of the Hamiltonian, and hence obtain the acceleration $a(r,v)$ due to the radiative force as a function of atom displacement from the MOT center, $r$, and velocity, $v$.
From $a(r,v)$, we can evolve atomic trajectories in phase space. These trajectories incorporate stochastic photon recoils from spontaneous emission as atoms repeatedly absorb and emit photons from the cooling laser. Further details of our numerical simulation analysis are given in the Supplementary Material (SM) [45].
IV Capture into Red-Detuned MOT
We consider a dual-frequency MOT as the capture MOT stage for Sn. The dual-frequency mechanism employs one circularly polarized laser beam red-detuned to the optical cycling transition, and a second, blue-detuned, co-propagating laser beam with orthogonal polarization (See Fig. 2). This generates an appreciable spatially restoring force on a Type-II transition [78, 88].
Figure 2: Dual frequency/polarization mechanism used for capture and conveyor belt MOTs of Sn atoms. Both configurations consist of two co-propagating laser beams with orthogonal circular polarizations. $\Delta$ is the detuning of a laser frequency from the $^{3}P_{1}\rightarrow{{}^{3}}P_{0}^{o}$ resonance.
The optimized and experimentally viable parameters for this stage are shown in Table 3. With Gaussian laser beams of $w_{\text{MOT}}=7$ mm ($1/e^{2}$ intensity radius), the scheme requires a power $P\sim\!200\,\text{mW}$ per beam of the MOT. This can be achieved with a total of $P\approx 200\,\text{mW}$ by using a single-beam configuration as in Ref. [36, 86]. The axial magnetic (B-)field gradient is $\partial B/\partial z=50\,\text{G/cm}$.
Table 3: Parameters used for Sn capture and compressed MOT simulations (see Fig. 3 and Fig. 6). For the compression stage, the parameters shown are those for the last stage of compression. Laser power $P_{i}$ and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ in the dual frequency/polarization scheme. $\partial B/\partial z$ is the axial B-field gradient. The Gaussian laser beam size ($1/e^{2}$ intensity radius) is $w_{\text{MOT}}=7$ $m$ in both MOT stages. In the results, $v_{\rm cap}$ is the capture velocity; $T$ and $\sigma$ are the temperature and cloud size; $P_{\rm exc}$ is the average population in the excited state.
MOT Stage $\partial B/\partial z$ (G/cm) Laser Parameters Results
$P_{i}$ (mW) $\Delta_{i}$ ($\Gamma$) $\hat{p}_{i}$ $v_{\rm cap}$ (m/s) $T$ (mK) $\sigma$ ($m$) $P_{\rm exc}$ (%)
Capture 50 172 1.0-1.0 $\sigma^{-}$ 28.5 225 1.9 7.8
34 +0.5+0.5 $\sigma^{+}$
Compress 150 5 0.5-0.5 $\sigma^{-}$ 8 0.5 0.6
1 +0.25+0.25 $\sigma^{+}$
A map of the acceleration experienced by Sn atoms in the capture MOT is shown in Fig. 3. The plot clearly indicates the presence of strong Doppler cooling at large velocities, for example with peak deceleration $a\approx-40$ mm/ms2 at $v\approx 20$ m/s and $r\approx-4$ mm. At lower velocities, $v\lesssim 5\,\text{m/s}$, the acceleration changes sign, indicating the presence of sub-Doppler heating. This behavior is typical for Type-II transitions [88], and can be more clearly seen in plots of the spatial and velocity-averaged forces derived from integrating $a(r,v)$ over $v$ and $r$, respectively, as shown in the SM  [45].
Figure 3: Acceleration experienced by Sn atoms in the capture MOT as a function of velocity and displacement towards the MOT center, orthogonal to the MOT’s magnetic quadrupole field symmetry axis, for the parameters listed in Table 3. Note the strong sub-Doppler heating at low velocities near the trap center, as seen from the reversed sign of acceleration in this region. This leads to a high equilibrium temperature in the MOT.
To determine capture velocity, we initialize atoms at r=14r=-14 mm, with increasing velocities $v$. An atom is deemed captured if, after $t=50$ ms of evolution, its trajectory has converged onto a steady-state oscillation in phase space, with the spatial extent of that oscillation $r_{\rm max}<w_{\rm MOT}$. We find that this condition is met for all trajectories for which the velocity reverses sign before the displacement along $\hat{r}$ exceeds $w_{\rm{MOT}}=7$ mm. We find a capture velocity of $v_{c}\approx 28.5$ m/s. This value is 2-3 times higher than what is experimentally measured for laser-cooled molecules like CaF and SrF [79, 89], due to the larger natural linewidth of the cooling transition and $\sim\!2\times$ higher photon momentum in Sn. Plots of trajectories for determining $v_{c}$ are shown in the SM [45].
To characterize the temperature and size of the MOT, we extend the trajectory of captured and trapped atoms by another $t=50$ ms to allow MOT observables to reach equilibrium. Using these trajectories, we calculate the rms displacement $\sigma$ (which we denote the cloud size), velocity $v_{\sigma}$, and temperature $T=mv_{\sigma}^{2}/k_{B}$, by averaging the position and velocity over the last 10 ms. The final size and temperature are averaged over ten such trajectories. For the capture MOT stage, we find $T\sim\!225\,\text{mK}$ and $\sigma\sim\!1.9\,\text{mm}$. For context, the Doppler temperature for this cooling transition in Sn is $T_{D}=760\,\mu\text{K}$. We attribute the extremely hot and large size of the capture red MOT to the large natural linewidth and Type-II nature of the optical cycling transition. A typical in-MOT trajectory is shown in the SM [45].
We note the possibility of two-photon ionization (TPI) when using the $^{3}P_{1}\rightarrow{{}^{3}}P_{0}^{o}$ optical cycling transition in Sn. This would manifest as one-body loss of the form $dN/dt=-\alpha N$. The loss rate due to TPI, $\alpha$, is given by:
$\alpha=\frac{\sigma P_{\text{exc}}I_{\text{tot}}}{\hbar\omega},$ (1)
where $\sigma$ is the TPI cross section, $P_{\text{exc}}$ is average population in the excited state, $I_{\text{tot}}$ is total laser intensity from all six passes, and $\omega$ is the angular frequency of the cycling transition. According to Ref. [90], Sn was calculated to have $\sigma\sim\!6\times 10^{-19}\,\text{cm}^{2}$ for this transition. Using the parameters in Table 3, we find $\alpha\sim\!0.1\,\text{s}^{-1}$, leading to a lifetime $\tau\sim\!10$ s in the capture MOT. For applications where the cold atoms are subsequently loaded into a far-detuned optical trap, this lifetime is not a limiting factor.
V White Light Slowing
For producing atoms with sufficiently low velocity to be captured in the MOT, we consider here the method of white light slowing (WLS) applied to the CBGB beam of atomic Sn. In WLS, the slowing laser spectrum is frequency-broadened to simultaneously cover many velocity classes. WLS is used in many molecular laser cooling experiments, enabling trapping of up to $\sim\!10^{4}$ molecules in a MOT [86, 32].
In this section, we use simulations to study WLS of a CBGB of Sn from the source region to the MOT loading region. The slowing laser detuning $\Delta_{s}$ is set roughly to resonance with the Doppler-shifted line of atoms at the center of the velocity range we wish to address. The laser spectrum is broadened using electro-optic modulators (EOMs). The sideband frequency $\Omega\sim\!1\Gamma$ is chosen to address as broad a range of velocities as possible. Commercial resonant-frequency EOMs for UV laser wavelengths can provide only modest modulation depth $\beta$, so we set $\beta\leq 1.5\,\text{rad}$ in our WLS simulation. To achieve sufficient spectral width with such limited values of $\beta$, we send the slowing laser through two successive EOMs with different frequencies and modulation depths. Finally, we use a converging slowing laser beam, which provides strongest deceleration near the CBGB source and a small transverse confining force, both desirable for efficient delivery of slowed atoms to the MOT region [91].
[…]
The WLS parameters that we determined to provide effective slowing of Sn atoms for capture into the MOT are shown in Table 4. The spectral profile of slowing light is shown in Fig. 4, and a plot of the atomic deceleration is shown in the SM [45]. The short excited-state lifetime of the optical cycling transition in atomic Sn makes it possible to apply a high scattering force to the atoms. We find that a slowing distance of $z_{\text{MOT}}=500\,\text{mm}$ is sufficient for effective WLS. This is shorter than in typical molecular MOT setups [29, 30, 33], resulting in a larger solid angle subtended by the MOT region and hence greater slowing efficiency [74].
Table 4: Slowing laser parameters used for simulations of white light slowing of Sn atoms. We assume a converging slowing beam with $w=7.5$ mm ($1/e^{2}$ intensity radius) at the MOT center ($z=500$ mm), half-angle $\theta^{\prime}=4.5$ mrad, and linear polarization. $P$ indicates the slowing laser power, $\Delta_{s}$ the slowing laser carrier frequency detuning, $\Omega_{i}$ the modulation frequency of each EOM, and $\beta_{i}$ the modulation depth of each EOM.
P (mW) $\Delta_{s}$ ($\Gamma$) $\Omega_{1}$ ($\Gamma$) $\Omega_{2}$ ($\Gamma$) $\beta_{1}$ (rad) $\beta_{2}$ (rad)
300 11.75-11.75 1.5 3.4 1.5 1.5
Figure 4: Spectral profile of slowing light. The red lines show the laser spectrum. The blue curve shows a convolution of this spectrum with a Lorentzian of linewidth $\Gamma$. The shaded gray region shows the initial longitudinal velocity distribution of Sn atoms from the CBGB, with the scale chosen such that the velocity and Doppler-shifted atomic resonance frequency coincide. The laser parameters are given in Table 4.
To find the capture efficiency of our WLS scheme, we perform simulations of atom trajectories undergoing WLS. Collisions of the atoms with He buffer gas immediately after exiting the source cell significantly affect CBGB properties  [84]. To account for this, we follow the procedure used in Ref. [74] to obtain an initial position distribution for the atoms: we convolve an initial uniform disk distribution (radius $R=1.5$ mm, the source cell aperture size) with a Gaussian distribution of FWHM $\Delta v_{T}\times z_{Q}/v_{L}$, where $z_{Q}=7.5$ mm is the zone of freezing (after which collisions between atoms and He are unlikely). We draw the initial transverse position of atoms at $z=z_{Q}$ from this convolved distribution. We then take the initial velocity distributions of Sn atoms at $z=z_{Q}$ to be the same as those of the SrF molecular laser cooling experiment (see Sec. II.3). Using these initial conditions, we run $1\times 10^{8}$ trajectory simulations. We solve the differential equations $\ddot{z}=a_{z}(\dot{z},r,z)$, $\ddot{x}=a_{\perp}x/r$, and $\ddot{y}=a_{\perp}y/r$, where the dot indicates a time derivative and $a_{z}$ is obtained using the parameters in Table 4, to propagate the atom trajectory from $z=z_{Q}$ to MOT capture region. We assume that an atom is captured by the MOT if it simultaneously satisfies two conditions [74]:
  1. 1.
    $v_{z}<v_{\text{cap}}$ at the location of the MOT ($z=z_{\text{MOT}}$).
  2. 2.
    $r<w_{\text{MOT}}$ at the location of the MOT. In other words, we take the MOT capture radius to be defined by the MOT laser beam radius.
Figure 5: Monte Carlo simulation of Sn atom trajectories propagating from the CBGB source to the MOT region under the influence of WLS, with parameters as described in the text. Only the longitudinal direction is shown. Trajectories in green (only every 20th trajectory is shown) indicate successful capture. Trajectories in red (only every 80000th trajectory is shown) indicate failure to capture. The computed efficiency of capture is $\sim\!1\times 10^{-4}$.
The result of this trajectory simulation is shown in Fig. 5. We find that the fraction of atoms captured is $\sim\!1\times 10^{-4}$. With our assumed Sn CBGB brightness and for the case of 120Sn (the most abundant isotope), this implies capture of $\sim\!10^{7}$ Sn atoms in the MOT. Note that similar efficiency calculations were performed in Ref. [74] for WLS of SrF molecules, where a capture efficiency over $100\times$ smaller was estimated.
VI Further Cooling and Trapping
VI.1 Compressed MOT
The parameters used for the Sn capture red MOT were chosen to maximize the capture velocity. This requires high scattering rate and leads to high temperature and large size, due to use of a Type-II cycling transition. To further cool the captured atoms, we propose to apply a compressed MOT (cMOT) stage after the capture MOT. This enables efficient transfer of cold atoms into subsequent cooling stages for direct loading of a conservative trap.
The cMOT is implemented by reducing the intensity and increasing the B-field gradient. The reduced scattering rate and higher B-field gradient drastically lowers the equilibrium temperature and tightens the spatial extent of the MOT [92]. We use our simulations to find experimentally viable parameters for the cMOT sequence (see Table 3). The compression sequence is shown in Fig. 6.
We find that decreasing the laser intensity to $3\%$ of its value in the capture MOT, tripling the axial B-field gradient $\partial B/\partial z$, and reducing the laser detunings $\Delta_{i}$ of both polarization components leads to a Sn cMOT with temperature $T\sim\!10\,\text{mK}$ and size $\sigma\sim\!500\,\mu\text{m}$. The average excited state population decreases from $\sim 8\%$ to $\sim 1\%$ during the compression sequence, leading to reduced heating from photon scattering. In total, the phase space density (PSD) increases by nearly five orders in magnitude from the capture MOT to the cMOT.
Figure 6: Simulated compression sequence for the Sn red MOT. Arrows indicate the progression of the sequence. The temperature $T$ and cloud size $\sigma$ are reduced by factors of 30 and 4, respectively, as laser power $P$ is reduced, laser detuning is decreased, and B-field gradient is increased. We anticipate a compression sequence time of $\sim\!25$ ms, commensurate with what is typically used in MOT experiments [86].
VI.2 Conveyor Belt Blue-Detuned MOT
We next consider further cooling of the Sn cloud using the “conveyor belt” blue-detuned MOT (CB MOT) technique [93]. A conventional blue MOT provides sub-Doppler cooling while maintaining spatial confinement in systems with Type-II transitions [94]. First demonstrated on the D2 line in 87Rb atoms [94], the blue MOT has also been successfully applied to a variety of laser-coolable molecules [95, 86, 96, 97].
The CB MOT is a novel variant of the blue MOT scheme that was recently theoretically proposed [93] and experimentally demonstrated in laser-cooled CaOH, CaF, BaF, and SrOH molecules [98, 97, 99, 39]. This new scheme provides an additional position-dependent trapping force, which enables formation of a more compressed atomic cloud compared to a conventional blue MOT. The CB MOT uses two blue-detuned laser beams with orthogonal polarizations and closely spaced ($1\Gamma$) frequencies (see Fig. 2).
[…]
Table 5: Parameters used for Sn CB MOT simulations (See Fig. 7). Laser power $P_{i}$ and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ in the dual frequency scheme. $\partial B/\partial z$ is the axial B-field gradient. The Gaussian laser beam size ($1/e^{2}$ intensity radius) is $w_{\text{MOT}}=7$ $m$ in both stages. $t_{D}$ is the compression time; $T$ and $\sigma$ are the temperature and cloud size.
Label $\partial B/\partial z$ (G/cm) Laser Parameters Results
$P_{i}$ (mW) $\Delta_{i}(\Gamma)$ $\hat{p}_{i}$ $t_{D}$ (ms) $T$ ($\mu\text{K}$) $\sigma$ ($\mu\text{m}$)
Stage 1 30 172 +3.2+3.2 $\sigma^{-}$ 10 75 20
172 +3.0+3.0 $\sigma^{+}$
Stage 2 40 114 +4.2+4.2 $\sigma^{-}$ 15 15 10
114 +4.0+4.0 $\sigma^{+}$
For the CB MOT of Sn atoms, we found a two-stage scheme to be effective. First, we use a relatively high laser intensity and low B-field gradient to maximize the number of atoms transferred from the cMOT to the CB MOT. Next, we reduce the laser intensity and increase the B-field gradient and laser detuning to further cool and compress the atomic cloud. We initialize atomic positions and velocities by randomly sampling from a distribution corresponding to the final size and temperature after the cMOT stage. Using the same method introduced in Section IV, we then simulate the atomic compression trajectories over a 50 ms duration. The relevant experimental parameters used in each CB MOT stage are shown in Table 5.
Fig. 7 shows atomic trajectories during each stage of the process. At the end of the first stage, which lasts $t\sim\!10\,\text{ms}$, we find an equilibrium temperature $T\sim\!75\,\mu\text{K}$ and cloud size $\sigma\sim\!20\,\mu\text{m}$. During the second stage, which lasts an additional $t\sim\!15\,\text{ms}$, the cloud is further cooled and compressed to $T\sim\!15\,\mu\text{K}$ and $\sigma\sim\!10\,\mu\text{m}$. In a real experiment, the atomic cloud will likely be hotter and larger than what was found in this simulation, due to our use of a semi-classical model and effects we neglected, like photon re-scattering and experimental imperfections. However, we do not expect the deviation of simulation from reality to be substantial [97].
VII Outlook
VII.1 Proposed Experimental Setup
Here, we briefly discuss a proposed experimental setup for realizing a Sn MOT, synthesizing our numerical simulation results with realistic experimental constraints. Commercially available laser systems at $\lambda=303\,\text{nm}$ can provide up to $P\approx 1\,\text{W}$ of power. We anticipate using acousto-optic modulators (AOM) to generate the laser frequencies needed for WLS and all MOT stages. Typical UV AOMs in a double-pass configuration have peak diffraction efficiency of $\sim\!45\%$ due to the strong polarization preference of UV AOM crystals.
In laser cooling of SrF molecules, WLS and capture MOT light are typically simultaneously on for optimal capture into the MOT. Since our results indicate $P\sim\!300\,\text{mW}$ is needed for efficient WLS of Sn atoms, we restrict the capture MOT power budget to $P\sim\!300\,\text{mW}$. Once capture is complete, WLS light can be shut off and used for subsequent MOT stages. Therefore, we restrict the CB MOT power budget to $P\sim\!450\,\text{mW}$. All told, this means the experiment will require a single-beam-pass MOT, where laser power is recycled through all six passes of the chamber. While not ideal for beam pointing stability and alignment, this type of beam delivery was routinely and successfully used in older-generation molecular laser cooling experiments for similar reasons [36, 86].
VII.2 Prospects for Further Cooling
[…]
VII.3 Summary
We have presented a realistic experimental scheme for laser cooling and trapping of Sn atoms. Based on the closely analogous structure of other Group IV elements (in particular, Si, Ge, and Pb), similar schemes should be effective in cooling and trapping these species as well. In forthcoming papers, we will explore the use of laser-cooled Group IV atoms for precision measurements and tests of fundamental physics, including atomic parity violation.
Acknowledgements
We thank Silviu M. Udrescu for discussions about the applications of laser-cooled tin atoms for precision measurement. G.Z. and Q.W. acknowledge support from the Air Force Office of Scientific Research. M.V. acknowledges support from the Natural Sciences and Engineering Research Council of Canada.
Supplemental Material: Simulated Laser Cooling and Magneto-Optical Trapping of Group IV Atoms
SM I Calculating the Metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ State Lifetime
The ground state of the proposed optical cycling transition for Group IV elements is the metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ state. This state only has one decay path, which goes to the absolute ground state ($s^{2}p^{2}\,{{}^{3}}P_{1}\leftrightarrow s^{2}p^{2}\,{{}^{3}}P_{0}$) via a pure M1 transition. Here, we analytically estimate the natural lifetime of the $s^{2}p^{2}\,{{}^{3}}P_{1}$ state to verify that the lifetime is long compared to typical magneto-optical trap (MOT) timescales. Throughout, we work in the $LS$-basis, which is approximately good for the lighter Group IV elements (C, Si, Ge) but requires some correction for the heavier Group IV elements (Sn, Pb). We compare our estimate to theoretical calculations done for C, Si, Sn, and Pb.
The magnetic dipole (M1) operator is given by:
$\hat{\mu}=-\frac{\mu_{B}}{\hbar}(g_{L}\mathbf{L}+g_{S}\mathbf{S}).$ (S1)
[…]
$\Gamma_{\rm M1}=\frac{\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}}\frac{|\langle g\|\hat{\mu}\|e\rangle|^{2}}{2J^{\prime}+1}.$ (S2)
We use standard results from angular momentum algebra and find:
$\langle L,S,J\|\mathbf{L}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{S,S^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}\langle L\|\mathbf{L}\|L^{\prime}\rangle,$ (S3)
and
$\langle L,S,J\|\mathbf{S}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{L,L^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\langle S\|\mathbf{S}\|S^{\prime}\rangle.$ (S4)
Using $\langle L\|\mathbf{L}\|L^{\prime}\rangle=\delta_{L,L^{\prime}}\sqrt{L(L+1)(2L+1)}$, and likewise for $\mathbf{S}$, we substitute and obtain:
$\Gamma_{\mathrm{M1}}$ $=\frac{\mu_{B}^{2}\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}}\frac{1}{2J^{\prime}+1}\bigg{(}\delta_{S,S^{\prime}}\delta_{L,L^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}$
$\quad\times\sqrt{L(L+1)(2L+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}+2\delta_{L,L^{\prime}}\delta_{S,S^{\prime}}(-1)^{L+S^{\prime}+J+1}$
$\quad\times\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{S(S+1)(2S+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\bigg{)}^{2}.$ (S5)
[…]
Table S1: Parameters of interest for the M1 transition $s^{2}p^{2}\,{{}^{3}}P_{1}\leftrightarrow s^{2}p^{2}\,{{}^{3}}P_{0}$ for all Group IV elements. $k$ denotes transition wavenumber, $\omega_{0}$ denotes transition energy, $\Gamma_{\rm M1}$ denotes natural linewidth, and $\tau_{g}$ denotes lifetime of the metastable state.
Element $k$ (cm-1) $\omega_{0}/2\pi$ (THz) $\Gamma_{\mathrm{M1}}/2\pi$ (Hz) $\tau_{g}$ (s)
Carbon (C) 16.42 0.49 $1.25\times 10^{-8}$ $1.27\times 10^{7}$
Silicon (Si) 77.12 2.31 $1.31\times 10^{-6}$ $1.22\times 10^{5}$
Germanium (Ge) 557.13 16.70 $4.95\times 10^{-4}$ 322
Tin (Sn) 1691.81 50.72 0.014 11.48
Lead (Pb) 7819.26 234.42 1.145 0.14
SM II Further Details on Atomic Tin MOT Simulations
Numerical simulations of all Sn MOT stages used a geometry where the $z$-axis is defined by the direction of the symmetry axis of the quadrupole magnetic field for the MOT. MOT laser beams in the $xy$-plane (which define the $x$ and $y$ axes) are oriented 45 degrees with respect to the atomic beam, which defines the axis $\hat{r}=(\hat{x}+\hat{y})/\sqrt{2}$. We use the same methodology as in Ref. [74] to analyze the behavior of the capture MOT, i.e. we only consider motion $\mathbf{v}\parallel\hat{r}$ and $\mathbf{d}\parallel\hat{r}$, where $\mathbf{v}$ and $\mathbf{d}$ denote 3-D velocity and displacement of atoms in the MOT. After solving the optical Bloch equations to obtain $a(r,v)$, we take one-dimensional cuts to elucidate the behavior of $a(r)$ and $a(v)$. These curves are defined by:
$a(r)=\frac{1}{2v_{\text{max}}}\int_{-v_{\text{max}}}^{v_{\text{max}}}a(r,v)\,dv,$ (S6)
and:
$a(v)=\frac{1}{2r_{\text{max}}}\int_{-r_{\text{max}}}^{r_{\text{max}}}a(r,v)\,dr.$ (S7)
Here, $v_{\text{max}}$ and $r_{\text{max}}$ are the maximum velocity and displacement of atoms in the capture MOT, which we find to be $v_{\text{max}}\approx 8$ m/s and $r_{\text{max}}\approx 3$ mm. Plots of these cuts are shown in Fig. S1(a) and (b), indicating the trapping and cooling behavior of the capture MOT stage. The sign reversal in the $a(v)$ curve at low velocities is a hallmark of sub-Doppler heating, a signature behavior of Type-II MOTs.
Fig. S1(c) shows how the capture velocity $v_{c}$ is determined, as described in the main text. We find that for velocities up to $v_{c}\sim 28.5$ m/s, atoms entering the MOT are slow enough to be captured. Captured atoms undergo a periodic, circular-like trajectory in phase-space, centered about the origin. This is indicative of sub-Doppler heating while trapped, due to the Type-II optical cycling transition.
Fig. S1(d) shows an example phase-space plot illustrative of how the temperature and size of the simulated capture MOT are determined. An atom is initialized at the phase-space origin and its trajectory propagated using the calculated acceleration $a(r,v)$. To account for random photon kicks from spontaneous emission during optical cycling, we give the atom a momentum kick in a random direction during each scattering event, which occurs with a rate $\sim P_{\rm exc}\Gamma$. The in-MOT trajectory with photon scattering is propagated for 50 ms to allow all observables to reach equilibrium. The temperature and size for each trajectory are calculated by computing the rms velocity and displacement in the last 10 ms of evolution. We average these values over 10 independent, stochastic trajectories to obtain the reported values for temperature $T$ and cloud size $\sigma$.
Figure S1: Analysis of the simulated Sn capture MOT. (a) Acceleration $a$ vs displacement $r$. The curve shows positional restoring behavior to the center of the trap. (b) Acceleration $a$ vs velocity $v$. The curve shows strong Doppler cooling leading to velocity damping for high velocities, as well as strong sub-Doppler heating at low velocities. (c) Phase space trajectory of Sn atoms entering the MOT region from the source beam. The arrowheads indicate the direction of time. Capture is possible for velocities up to $v_{c}\sim\!28.5$ m/s. (d) In-MOT trajectory of Sn atoms in the capture MOT, accounting for spontaneous photon emission due to optical cycling. The arrowhead indicates the direction of time. By computing the rms velocity and displacement averaged over the last 10 ms of the trajectory, we find the MOT temperature $T\sim\!225$ mK and cloud size $\sigma\sim\!1.9$ $m$.
Compressed MOT simulations were analyzed using the same protocol as for the capture MOT, since the two stages are identical except for the reduced laser power, reduced laser detuning, and increased $B$-field gradient of the compression stage compared to the capture stage. The main difference in analysis is that a finer grid for interpolation of $a(r,v)$ values is used in the compression stage, to account for the fact that the MOT temperature and size reduces during this sequence. Conveyor belt blue MOT simulations were also analyzed similarly, with emphasis placed on the rms positional compression of the MOT cloud during this stage.
SM III Further Details on Tin White Light Slowing Simulations
Numerical simulations of white light slowing (WLS) of Sn used a geometry where the atomic beam defines the $z$-axis. The slowing laser polarization is linear and defines the $x$-axis. Sn atoms exit the CBGB source with longitudinal velocity $v_{z}$ and transverse velocity $v_{\perp}$ drawn from independent Gaussian distributions, as described in the main text. Their transverse displacement $r$ is drawn from a probability distribution which is the convolution of a Gaussian of FWHM $\approx 4$ mm and a 2-D uniform disk of radius $R=1.5$ mm centered on the $z$-axis. Sn atoms are initialized at $z_{Q}=7.5$ mm, which defines the end of the “zone of freezing” where re-thermalizing collisions of Sn atoms with He buffer gas cease.
We consider realistic experimental parameters in our simulations of white light slowing. Spectral broadening at ultraviolet laser wavelengths can be achieved using resonant crystal electro-optic modulators (EOMs), whose modulation depths typically cannot exceed $\beta\sim 2$ rad. For sufficient laser broadening, we use a dual EOM configuration where the slowing laser passes through two EOMs in succession ($\beta_{1}=\beta_{2}=1.5$ rad, $\Omega_{1}=1.5\Gamma$, $\Omega_{2}=3.4\Gamma$). The broadened laser spectrum can be analytically calculated using the Jacobi-Anger expansion. The time dependence of the laser electromagnetic field is given by:
$e^{i\omega t+i\beta_{1}\sin(\Omega_{1}t)+i\beta_{2}\sin(\Omega_{2}t)}=e^{i\omega t}\left(J_{0}(\beta_{1})+\sum_{k=1}^{\infty}J_{k}(\beta_{1})e^{ik\Omega_{1}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{1})e^{-ik\Omega_{1}t}\right)\times\\ \left(J_{0}(\beta_{2})+\sum_{k=1}^{\infty}J_{k}(\beta_{2})e^{ik\Omega_{2}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{2})e^{-ik\Omega_{2}t}\right),$ (S8)
where $J_{k}(\beta)$ is the $k$th order Bessel function of the first kind. This corresponds to having spectral power in the original carrier $\omega$ as well as at frequency sidebands $\omega\pm k\Omega_{1}\pm m\Omega_{2}$ where $k,m$ are integers. The powers in each spectral component are determined by the values of $J_{k}(\beta_{1})$ and $J_{m}(\beta_{2})$.
After applying this broadened laser spectrum in the WLS simulations, we solve the OBEs and obtain the longitudinal deceleration $a_{z}(v_{z},r)$. We incorporate a converging slowing beam with half-angle $\theta^{\prime}=4.5$ mrad by solving the OBEs for $a_{z}(v_{z},r)$ as a function of longitudinally-varying laser intensity $s(z)$. Additionally, we take the transverse deceleration due to the converging slowing beam to be $a_{\perp}=a_{z}\tan\theta^{\prime}$. Interpolation across $z$ values from $z=0$ mm to $z=500$ mm enables us to obtain deceleration curves $a_{z}(v_{z},r,z)$, which are shown in Fig. S2. The calculated $a_{z}(v_{z},r,z)$ and $a_{\perp}(v_{z},r,z)$ are used to solve the differential equations for the trajectory of Sn atoms as they are slowed from source to MOT region.
Figure S2: Three-dimensional plot of the longitudinal deceleration $a_{z}(v_{z},z)$ imparted on Sn atoms due to white light slowing. Profiles for three different transverse displacements from the atomic beam axis $\hat{z}$ are shown: $r=0$ mm (blue), $r=3$ mm (green), and $r=6$ mm (brown). These show the decrease in slowing force as $r$ increases. The longitudinal intensity gradient due to the converging slowing beam is apparent as the peak value of $a_{z}$ gets increasingly negative for decreasing $z$.

### v2

@@
 Group 14 or tetrel) atoms silicon (Si), germanium (Ge), tin (Sn), and lead (Pb).
-These elements each possess a strong Type-II transition ($J\rightarrow J^{\prime}=J-1$) between the metastable $s^{2}p^{2}\,{}^{3}P_{1}$ state and the excited $s^{2}ps^{\prime}\,{}^{3}P_{0}^{o}$ state at an accessible laser wavelength, making them amenable to laser cooling and trapping.
+These elements each possess a strong Type-II transition ($J\rightarrow J^{\prime}=J-1$) between the metastable $s^{2}p^{2}\,{}^{3}\!P_{1}$ state and the excited $s^{2}ps^{\prime}\,{}^{3}\!P_{0}^{\circ}$ state at an accessible laser wavelength, making them amenable to laser cooling and trapping.
 We focus on the application of this scheme to Sn, which has several features that make it attractive for precision measurement applications.
@@
 Adding a new class of elements to the list of laser-coolable species would constitute another leap in the quest to control and manipulate quantum matter.
-In this work, we present numerical simulations to verify a scheme for laser cooling and trapping of Group IV elements — silicon (Si), germanium (Ge), tin (Sn), and lead (Pb) — using a Type-II transition.
+In this work, we present numerical simulations to verify a scheme for laser cooling and trapping of nearly all [^1] The vacuum UV transition wavelength for carbon (C) cannot currently be produced with sufficient continuous-wave power for laser cooling and trapping (see Table 1), so we do not consider it here.
+Group IV elements — silicon (Si), germanium (Ge), tin (Sn), and lead (Pb) — using a Type-II transition.
 Among these elements, previous attempts have been made to study pathways towards producing ultracold gases of Si atoms [53, 32, 82, 87], but a Si MOT has not been realized to date.
@@
 In their ground state, all Group IV elements have four valence electrons in a $s^{2}p^{2}$ atomic orbital configuration.
-Generically, there are five even parity states in the ground-state manifold, with terms $^{3}P_{0,1,2},{{}^{1}D}_{2}$, and $^{1}S_{0}$ in order of energy from lowest to highest.
-The lowest-lying odd parity excited state has configuration $s^{2}ps^{\prime}$ and term $^{3}P_{0}^{o}$.
-[^1] C and Si also have a $sp^{3}$ $^{5}S_{2}^{o}$ state below the $s^{2}ps^{\prime}$ $^{3}P^{o}$ states and above the $s^{2}p^{2}$ $^{1}S_{0}$ state, but this does not affect our discussion.
-All elements in this group have multiple spin-zero ($I=0$) even isotopes and several spinful odd isotopes; in this paper we focus only on the even isotopes.
+This leads to five even parity states in the ground-state manifold, with terms $^{3}\!P_{0,1,2},{{}^{1}\!D}_{2}$, and $^{1}\!S_{0}$ in order of energy from lowest to highest.
+The four excited states in this manifold cannot decay via electric dipole (E1) radiation and hence are all long-lived.
+The lowest-lying odd parity excited state has configuration $s^{2}ps^{\prime}$ and term $^{3}\!P_{0}^{\circ}$.
+[^2] C and Si also have a $sp^{3}$ $^{5}\!S_{2}^{\circ}$ state below the $s^{2}ps^{\prime}$ $^{3}\!P^{\circ}$ states and above the $s^{2}p^{2}$ $^{1}\!S_{0}$ state, but this does not affect our discussion.
 A level diagram of this structure is shown in Fig.
@@
 […]
-We propose using the metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ to excited $s^{2}ps^{\prime}\,{{}^{3}}P_{0}^{o}$ transition in the Group IV elements as a pure Type-II optical cycling transition that is amenable to laser cooling and trapping.
+We propose using the metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ to excited $s^{2}ps^{\prime}\,{{}^{3}}\!P_{0}^{\circ}$ transition in the Group IV elements as a pure Type-II optical cycling transition that is amenable to laser cooling and trapping.
 This transition is dipole-allowed with large natural linewidth.
@@
 Table 1 lists the relevant properties of this transition across all Group IV elements, as well as the metastable state lifetime.
-In the case of Si, Ge, Sn, and Pb, the transition wavelength can be produced with current ultraviolet (UV) laser technology.
-In the case of C, the transition wavelength is too short to currently be produced for laser cooling and trapping, so we do not consider it further in this work.
-Table 1: Optical transition wavelength $\lambda_{c}$, natural linewidth $\Gamma$, and saturation intensity $I_{\text{sat}}$ of the $s^{2}p^{2}\,{{}^{3}}P_{1}\rightarrow s^{2}ps^{\prime}\,{{}^{3}}P_{0}^{o}$ transition proposed for use in laser cooling and trapping of Group IV elements [44].
+In the case of Si, Ge, Sn, and Pb, the transition wavelength can be produced with sufficient power using current UV laser technology.
+Table 1: Optical transition wavelength $\lambda_{c}$, natural linewidth $\Gamma$, and saturation intensity $I_{\text{sat}}$ of the $s^{2}p^{2}\,{{}^{3}}\!P_{1}\rightarrow s^{2}ps^{\prime}\,{{}^{3}}\!P_{0}^{\circ}$ transition proposed for use in laser cooling and trapping of Group IV elements [52].
 The metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ state lifetime $\tau_{g}$ is also listed [52].
-More details on calculating $\tau_{g}$ are given in the Supplemental Material (SM) [45].
-Element $\lambda_{c}$ (nm) $\Gamma$ (MHz) $I_{\text{sat}}$ (mW/cm2) $\tau_{g}$ (s)
-C 166 $2\pi\times 55$ 1572 $1.3\times 10^{7}$
-Si 252 $2\pi\times 35$ 286 $1.2\times 10^{5}$
-Ge 271 $2\pi\times 45$ 296 $3.2\times 10^{2}$
-Sn 303 $2\pi\times 32$ 149 11
-Pb 368 $2\pi\times 22$ 58 0.14
-We note that a more conventional-seeming optical cycling transition is also present: $s^{2}p^{2}\,{}^{3}P_{2}$ to $sp^{3}\,{{}^{3}}D_{3}^{o}$ (for Si) or $s^{2}pd\,{{}^{3}}D_{3}^{o}$ (for Ge and Sn), i.e.
-the excited state is the lowest-lying odd-parity $D$ state.
-This was used to attempt laser cooling of Si ($\lambda=222\,\text{nm}$) [42].
-This cycling transition is leaky, with an estimated branching ratio $^{3}D_{3}^{o}\rightarrow{{}^{1}}D_{2}$ of $\sim\!0.02\%-50\%$ depending on the element [44], necessitating a repumper laser to close the cycling transition.
-However, inclusion of this repumper creates a $\Lambda$-system, effectively making a Type-II cycling transition with two UV lasers on two different transitions.
-The cycling transition we propose is simpler and should lead to more effective laser cooling and trapping of Group IV elements.
-In addition to supporting a Type-II transition, all Group IV elements possess four linearly independent clock transitions between the long-lived $s^{2}p^{2}$ states.
-Details on these transitions, which can only be of magnetic dipole (M1) or electric quadrupole (E2) type, are listed in Table 2.
+Details on calculating $\tau_{g}$ are given in Appendix A.
+Element $\lambda_{c}$ (nm) $\Gamma/2\pi$ (MHz) $I_{\text{sat}}$ (mW/cm2) $\tau_{g}$ (s)
+C 166 $55$ 1600 $1\times 10^{7}$
+Si 252 $35$ 290 $1\times 10^{5}\!$
+Ge 271 $45$ 300 $3\times 10^{2}$
+Sn 303 $32$ 150 11
+Pb 368 $22$ 60 0.14
+II.2 Applications of Laser-Cooled Group IV Atoms
+Ultracold atomic gases of Group IV elements have potential applications ranging from improved quantum sensing and novel quantum computing platforms to precision tests of fundamental physics and searches for physics beyond the Standard Model (BSM).
+Initial efforts to produce an ultracold atomic gas of Si atoms focused on envisioned applications to solid-state-based nuclear spin quantum computing [50, 54] and quantum sensing [68].
+Resonant photoionization of ultracold radioactive $^{31}\text{Si}$ atoms in a MOT, followed by ion-guiding optics, would enable nm-scale precision implantation of $^{31}\text{Si}^{+}$ ions in a Si substrate [82, 41].
+Subsequent $\beta$-decay of these nuclei ($\tau\sim\!3$ hrs) would create an array of $^{31}\text{P}^{+}$ ions doped in the Si substrate, realizing the building blocks of the Kane quantum computing architecture [50].
+In addition to supporting a Type-II cycling transition, all Group IV elements possess various clock transitions that can address all four linearly independent energy differences between the long-lived $s^{2}p^{2}$ states.
+Details on these transitions, which can only be of magnetic dipole (M1) or electric quadrupole (E2) type, are listed in Table 2 for the case of Sn.
 Table 2: List of narrow linewidth clock transitions between $s^{2}p^{2}$ states in Group IV elements.
@@
 Atomic Sn
-State Transition Type $\lambda$ (nm) $\Gamma_{\rm p}$ (Hz) $\Gamma_{\rm tot}$ (Hz)
-$^{3}P_{1}$ $\leftrightarrow\,^{3}P_{0}$ M1 5911 $2\pi\times 0.01$ $2\pi\times 0.01$
-$^{3}P_{2}$ $\leftrightarrow\,^{3}P_{0}$ E2 2917 $2\pi\times 10^{-4}$ $2\pi\times 0.01$
-$\leftrightarrow\,^{3}P_{1}$ M1/E2 5761 $2\pi\times 0.01$
-$^{1}D_{2}$ $\leftrightarrow\,^{3}P_{0}$   E2* 1161 $2\pi\times 10^{-5}$ $2\pi\times 0.16$
-$\leftrightarrow\,^{3}P_{1}$   M1/E2* 1445 $2\pi\times 0.07$
-$\leftrightarrow\,^{3}P_{2}$   M1/E2* 1929 $2\pi\times 0.08$
-$^{1}S_{0}$ $\leftrightarrow\,^{3}P_{1}$   M1* 646 $2\pi\times 1.13$ $2\pi\times 1.42$
-$\leftrightarrow\,^{3}P_{2}$   E2* 728 $2\pi\times 0.11$
-$\leftrightarrow\,^{1}D_{2}$ E2 1170 $2\pi\times 0.18$
-II.2 Applications of Laser-Cooled Group IV Atoms
-Ultracold atomic gases of Group IV elements have potential applications ranging from improved quantum sensing and novel quantum computing platforms to precision tests of fundamental physics and searches for physics beyond the Standard Model (BSM).
-Initial efforts to produce an ultracold atomic gas of Si atoms focused on envisioned applications to solid-state-based nuclear spin quantum computing [47, 48] and quantum sensing [49].
-Resonant photoionization of ultracold radioactive $^{31}\text{Si}$ atoms in a MOT, followed by ion-guiding optics, would enable nm-scale precision implantation of $^{31}\text{Si}^{+}$ ions in a Si substrate [42, 50].
-Subsequent $\beta$-decay of these ions ($\tau\sim\!3$ hrs) would create an array of $^{31}\text{P}^{+}$ ions doped in the Si substrate, realizing the building blocks of the Kane quantum computing architecture [47].
-The plethora of accessible clock transitions in the ground $s^{2}p^{2}$ configuration can be utilized for a variety of precision measurement experiments.
-Sn is particularly attractive for such applications because it possesses a chain of seven stable, $I=0$ isotopes (the longest found in nature).
+State Transition Type $\lambda$ (nm) $\Gamma_{\text{p}}/2\pi$ (Hz) $\Gamma_{\text{tot}}/2\pi$ (Hz)
+$^{3}\!P_{1}$ $\leftrightarrow\,^{3}\!P_{0}$ M1 5911 $0.01$ $0.01$
+$^{3}\!P_{2}$ $\leftrightarrow\,^{3}\!P_{0}$ E2 2917 $10^{-4}$ $0.01$
+$\leftrightarrow\,^{3}\!P_{1}$ M1/E2 5761 $0.01$
+$^{1}\!D_{2}$ $\leftrightarrow\,^{3}\!P_{0}$   E2* 1161 $10^{-5}$ $0.16$
+$\leftrightarrow\,^{3}\!P_{1}$   M1/E2* 1445 $0.07$
+$\leftrightarrow\,^{3}\!P_{2}$   M1/E2* 1929 $0.08$
+$^{1}\!S_{0}$ $\leftrightarrow\,^{3}\!P_{1}$   M1* 646 $1.13$ $1.4$
+$\leftrightarrow\,^{3}\!P_{2}$   E2* 728 $0.11$
+$\leftrightarrow\,^{1}\!D_{2}$ E2 1170 $0.18$
+The plethora of metastable states in the ground $s^{2}p^{2}$ configuration provides several accessible clock transitions that can be utilized for a variety of precision measurement experiments.
+Sn is particularly attractive for such applications because it possesses a chain of seven stable $I=0$ isotopes (the longest found in nature).
 The combination of many clock transitions and many $I=0$ isotopes makes laser-cooled and trapped Sn atoms a promising platform for probing non-linearity in King plot isotope shift measurements [9, 36].
@@
 This unique class of molecule was recently proposed as a promising candidate for next-generation precision measurements of the electron electric dipole moment (eEDM) [90].
-II.3 Scheme for Laser Cooling and Trapping with the $^{3}P_{1}\rightarrow{{}^{3}P}_{0}^{o}$ Transition
+II.3 Scheme for Laser Cooling and Trapping with the $^{3}\!P_{1}\rightarrow{{}^{3}\!P}_{0}^{\circ}$ Transition
 […]
+We note, in passing, that a more conventional-seeming optical cycling transition is also present: $s^{2}p^{2}\,{}^{3}\!P_{2}\leftrightarrow sp^{3}\,{{}^{3}}\!D_{3}^{\circ}$ (for Si) or $s^{2}pd\,{{}^{3}}\!D_{3}^{\circ}$ (for Ge and Sn), i.e.
+the excited state is the lowest-lying odd-parity triplet state with $J=3$.
+This was used to attempt laser cooling of Si ($\lambda=222\,\text{nm}$) [82].
+This cycling transition is leaky, with an estimated branching ratio $^{3}\!D_{3}^{\circ}\rightarrow s^{2}p^{2}{{}^{1}}\!D_{2}$ of $\sim\!0.02\%-50\%$ depending on the element [52], necessitating a repumper laser to close the cycling transition.
+However, including this repumper creates a $\Lambda$-system, effectively making a Type-II cycling transition with two UV lasers on two different transitions.
+Our approach is simpler and should lead to more effective laser cooling and trapping of Group IV elements.
 The presence of dark Zeeman sublevels in the ground state necessitates a mechanism to destabilize such states during optical cycling, lest population accumulate there and cause cooling and trapping to cease [11].
@@
 The initial transverse velocity distribution is also roughly Gaussian, with mean $\bar{v}_{T}=0\,\text{m/s}$ and FWHM $\Delta{v_{T}}=75\,\text{m/s}$.
-We assume a beam brightness across all isotopes of $\mathcal{B}\sim\!1.5\times 10^{12}$ atoms in the $^{3}P_{1}$ metastable state/steradian/pulse, commensurate with the measured brightness of Sr atoms in the SrF beam [85] after accounting for the metastable state population.
+We assume a beam brightness across all isotopes of $\mathcal{B}\sim\!1.5\times 10^{12}$ atoms in the $^{3}\!P_{1}$ metastable state per steradian per pulse, commensurate with the measured brightness of Sr atoms in the SrF beam [49] after accounting for the metastable state population.
 We remark that using a thermal beam source of atoms may also be viable, but we do not explore this path here.
@@
 These trajectories incorporate stochastic photon recoils from spontaneous emission as atoms repeatedly absorb and emit photons from the cooling laser.
-Further details of our numerical simulation analysis are given in the Supplementary Material (SM) [45].
+Further details of our numerical simulation analysis are given in Appendix B.
 IV Capture into Red-Detuned MOT
@@
 Both configurations consist of two co-propagating laser beams with orthogonal circular polarizations.
-$\Delta$ is the detuning of a laser frequency from the $^{3}P_{1}\rightarrow{{}^{3}}P_{0}^{o}$ resonance.
+$\Delta$ is the detuning of a laser frequency from the $^{3}\!P_{1}\rightarrow{{}^{3}}\!P_{0}^{\circ}$ resonance.
 The optimized and experimentally viable parameters for this stage are shown in Table 3.
-With Gaussian laser beams of $w_{\text{MOT}}=7$ mm ($1/e^{2}$ intensity radius), the scheme requires a power $P\sim\!200\,\text{mW}$ per beam of the MOT.
-This can be achieved with a total of $P\approx 200\,\text{mW}$ by using a single-beam configuration as in Ref.
+With Gaussian laser beams of $w_{\text{MOT}}=7$ mm ($1/e^{2}$ intensity radius), the scheme requires a power $P_{\text{beam}}\sim\!200\,\text{mW}$ per beam of the MOT.
+This can be achieved with a total of $P_{\text{tot}}\approx 200\,\text{mW}$ by using a single-beam configuration as in Ref.
 [56, 48].
@@
 For the compression stage, the parameters shown are those for the last stage of compression.
-Laser power $P_{i}$ and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ in the dual frequency/polarization scheme.
+Laser power $P_{i}$, saturation parameter $s_{i}$, and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ (for the given polarization, we assume the beam wave vector $\hat{k}_{i}$ points towards $\hat{x}$ and that the B-field increases along $\hat{x}$) in the dual frequency/polarization scheme.
 $\partial B/\partial z$ is the axial B-field gradient.
@@
 In the results, $v_{\rm cap}$ is the capture velocity; $T$ and $\sigma$ are the temperature and cloud size; $P_{\rm exc}$ is the average population in the excited state.
-MOT Stage $\partial B/\partial z$ (G/cm) Laser Parameters Results
-$P_{i}$ (mW) $\Delta_{i}$ ($\Gamma$) $\hat{p}_{i}$ $v_{\rm cap}$ (m/s) $T$ (mK) $\sigma$ ($m$) $P_{\rm exc}$ (%)
-Capture 50 172 1.0-1.0 $\sigma^{-}$ 28.5 225 1.9 7.8
-34 +0.5+0.5 $\sigma^{+}$
-Compress 150 5 0.5-0.5 $\sigma^{-}$ 8 0.5 0.6
-1 +0.25+0.25 $\sigma^{+}$
+MOT Stage $\frac{\partial B}{\partial z}$ (G/cm) Laser Parameters Results
+$P_{i}$ (mW) $s_{i}$ $\Delta_{i}$ ($\Gamma$) $\hat{p}_{i}$ $v_{\rm cap}$ (m/s) $T$ (mK) $\sigma$ ($m$) $P_{\rm exc}$ (%)
+Capture 50 172 1.5 1.0-1.0 $\sigma^{-}$ 28.5 225 1.9 7.8
+34 0.3 +0.5+0.5 $\sigma^{+}$
+Compress 150 5 0.05 0.5-0.5 $\sigma^{-}$ 8 0.5 0.6
+1 0.01 +0.25+0.25 $\sigma^{+}$
 A map of the acceleration experienced by Sn atoms in the capture MOT is shown in Fig.
@@
 At lower velocities, $v\lesssim 5\,\text{m/s}$, the acceleration changes sign, indicating the presence of sub-Doppler heating.
-This behavior is typical for Type-II transitions [88], and can be more clearly seen in plots of the spatial and velocity-averaged forces derived from integrating $a(r,v)$ over $v$ and $r$, respectively, as shown in the SM  [45].
+This behavior is typical for Type-II transitions [27], and can be more clearly seen in plots of the spatial and velocity-averaged forces derived from integrating $a(r,v)$ over $v$ and $r$, respectively, as shown in Appendix B.
 Figure 3: Acceleration experienced by Sn atoms in the capture MOT as a function of velocity and displacement towards the MOT center, orthogonal to the MOT’s magnetic quadrupole field symmetry axis, for the parameters listed in Table 3.
@@
 This value is 2-3 times higher than what is experimentally measured for laser-cooled molecules like CaF and SrF [101, 89], due to the larger natural linewidth of the cooling transition and $\sim\!2\times$ higher photon momentum in Sn.
-Plots of trajectories for determining $v_{c}$ are shown in the SM [45].
+Plots of trajectories for determining $v_{c}$ are shown in Appendix B.
 To characterize the temperature and size of the MOT, we extend the trajectory of captured and trapped atoms by another $t=50$ ms to allow MOT observables to reach equilibrium.
@@
 For context, the Doppler temperature for this cooling transition in Sn is $T_{D}=760\,\mu\text{K}$.
-We attribute the extremely hot and large size of the capture red MOT to the large natural linewidth and Type-II nature of the optical cycling transition.
-A typical in-MOT trajectory is shown in the SM [45].
-We note the possibility of two-photon ionization (TPI) when using the $^{3}P_{1}\rightarrow{{}^{3}}P_{0}^{o}$ optical cycling transition in Sn.
+We attribute the extremely hot temperature and large size of the capture red MOT to the large natural linewidth and Type-II nature of the optical cycling transition; a similarly hot MOT was observed in Ref.
+[71].
+A typical in-MOT trajectory is shown in Appendix B.
+We note the possibility of two-photon ionization (TPI) when using the $^{3}\!P_{1}\rightarrow{{}^{3}}\!P_{0}^{\circ}$ optical cycling transition in Sn.
 This would manifest as one-body loss of the form $dN/dt=-\alpha N$.
@@
 The spectral profile of slowing light is shown in Fig.
-4, and a plot of the atomic deceleration is shown in the SM [45].
+4, and a plot of the atomic deceleration is shown in Appendix C.
 The short excited-state lifetime of the optical cycling transition in atomic Sn makes it possible to apply a high scattering force to the atoms.
@@
 We assume a converging slowing beam with $w=7.5$ mm ($1/e^{2}$ intensity radius) at the MOT center ($z=500$ mm), half-angle $\theta^{\prime}=4.5$ mrad, and linear polarization.
-$P$ indicates the slowing laser power, $\Delta_{s}$ the slowing laser carrier frequency detuning, $\Omega_{i}$ the modulation frequency of each EOM, and $\beta_{i}$ the modulation depth of each EOM.
-P (mW) $\Delta_{s}$ ($\Gamma$) $\Omega_{1}$ ($\Gamma$) $\Omega_{2}$ ($\Gamma$) $\beta_{1}$ (rad) $\beta_{2}$ (rad)
-300 11.75-11.75 1.5 3.4 1.5 1.5
+$P$ indicates the slowing laser power, $\bar{s}$ the saturation parameter spatially averaged along the slowing beamline axis, $\Delta_{s}$ the slowing laser carrier frequency detuning, $\Omega_{i}$ the modulation frequency of each EOM, and $\beta_{i}$ the modulation depth of each EOM.
+$P$ (mW) $\bar{s}$ $\Delta_{s}$ ($\Gamma$) $\Omega_{1}$ ($\Gamma$) $\Omega_{2}$ ($\Gamma$) $\beta_{1}$ (rad) $\beta_{2}$ (rad)
+300 3.3 11.75-11.75 1.5 3.4 1.5 1.5
 Figure 4: Spectral profile of slowing light.
@@
 Arrows indicate the progression of the sequence.
-The temperature $T$ and cloud size $\sigma$ are reduced by factors of 30 and 4, respectively, as laser power $P$ is reduced, laser detuning is decreased, and B-field gradient is increased.
+The temperature $T$ and cloud size $\sigma$ are reduced by factors of 30 and 4, respectively, as laser power $P$ is reduced ($P_{\text{capture}}$ is the laser power used for capture MOT stage), laser detuning is decreased, and B-field gradient is increased.
 We anticipate a compression sequence time of $\sim\!25$ ms, commensurate with what is typically used in MOT experiments [48].
@@
 7).
-Laser power $P_{i}$ and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ in the dual frequency scheme.
+Laser power $P_{i}$, saturation parameter $s_{i}$, and detuning $\Delta_{i}$ are given for each polarization component $\hat{p}_{i}$ (for the given polarization, we assume same beam wave vector $\hat{k}_{i}$ direction as in previous MOT stages, see Table 3) in the dual frequency scheme.
 $\partial B/\partial z$ is the axial B-field gradient.
@@
 $t_{D}$ is the compression time; $T$ and $\sigma$ are the temperature and cloud size.
-Label $\partial B/\partial z$ (G/cm) Laser Parameters Results
-$P_{i}$ (mW) $\Delta_{i}(\Gamma)$ $\hat{p}_{i}$ $t_{D}$ (ms) $T$ ($\mu\text{K}$) $\sigma$ ($\mu\text{m}$)
-Stage 1 30 172 +3.2+3.2 $\sigma^{-}$ 10 75 20
-172 +3.0+3.0 $\sigma^{+}$
-Stage 2 40 114 +4.2+4.2 $\sigma^{-}$ 15 15 10
-114 +4.0+4.0 $\sigma^{+}$
+Label $\frac{\partial B}{\partial z}$ (G/cm) Laser Parameters Results
+$P_{i}$ (mW) $s_{i}$ $\Delta_{i}(\Gamma)$ $\hat{p}_{i}$ $t_{D}$ (ms) $T$ ($\mu\text{K}$) $\sigma$ ($\mu\text{m}$)
+Stage 1 30 172 1.5 +3.2+3.2 $\sigma^{-}$ 10 75 20
+172 1.5 +3.0+3.0 $\sigma^{+}$
+Stage 2 40 114 1.0 +4.2+4.2 $\sigma^{-}$ 15 15 10
+114 1.0 +4.0+4.0 $\sigma^{+}$
 For the CB MOT of Sn atoms, we found a two-stage scheme to be effective.
@@
 […]
-VII.3 Summary
+VIII Summary
 We have presented a realistic experimental scheme for laser cooling and trapping of Sn atoms.
@@
 In forthcoming papers, we will explore the use of laser-cooled Group IV atoms for precision measurements and tests of fundamental physics, including atomic parity violation.
-Acknowledgements
+Acknowledgements.
 We thank Silviu M.
@@
 and Q.W.
-acknowledge support from the Air Force Office of Scientific Research.
+acknowledge support from the Air Force Office of Scientific Research (AFOSR MURI Grant No.
+FA9550-21-1-0069).
 M.V.
-acknowledges support from the Natural Sciences and Engineering Research Council of Canada.
-Supplemental Material: Simulated Laser Cooling and Magneto-Optical Trapping of Group IV Atoms
-SM I Calculating the Metastable $s^{2}p^{2}\,{{}^{3}}P_{1}$ State Lifetime
+acknowledges support from the Natural Sciences and Engineering Research Council of Canada (NSERC).
+Data Availability
+The data that support the findings of this article are openly available [107].
+Appendix A Calculating the Metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ State Lifetime
 The ground state of the proposed optical cycling transition for Group IV elements is the metastable $s^{2}p^{2}\,{{}^{3}}\!P_{1}$ state.
@@
 The magnetic dipole (M1) operator is given by:
-$\hat{\mu}=-\frac{\mu_{B}}{\hbar}(g_{L}\mathbf{L}+g_{S}\mathbf{S}).$ (S1)
+$\hat{\mu}=-\frac{\mu_{B}}{\hbar}(g_{L}\mathbf{L}+g_{S}\mathbf{S}).$ (2)
 […]
-$\Gamma_{\rm M1}=\frac{\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}}\frac{|\langle g\|\hat{\mu}\|e\rangle|^{2}}{2J^{\prime}+1}.$ (S2)
+$\Gamma_{\rm M1}=\frac{\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}\!}\frac{|\langle g\|\hat{\mu}\|e\rangle|^{2}}{2J^{\prime}+1}.$ (3)
 We use standard results from angular momentum algebra and find:
-$\langle L,S,J\|\mathbf{L}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{S,S^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}\langle L\|\mathbf{L}\|L^{\prime}\rangle,$ (S3)
+$\langle L,S,J\|\mathbf{L}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{S,S^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}\langle L\|\mathbf{L}\|L^{\prime}\rangle,$ (4)
 and
-$\langle L,S,J\|\mathbf{S}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{L,L^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\langle S\|\mathbf{S}\|S^{\prime}\rangle.$ (S4)
+$\langle L,S,J\|\mathbf{S}\|L^{\prime},S^{\prime},J^{\prime}\rangle=\delta_{L,L^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\langle S\|\mathbf{S}\|S^{\prime}\rangle.$ (5)
 Using $\langle L\|\mathbf{L}\|L^{\prime}\rangle=\delta_{L,L^{\prime}}\sqrt{L(L+1)(2L+1)}$, and likewise for $\mathbf{S}$, we substitute and obtain:
-$\Gamma_{\mathrm{M1}}$ $=\frac{\mu_{B}^{2}\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}}\frac{1}{2J^{\prime}+1}\bigg{(}\delta_{S,S^{\prime}}\delta_{L,L^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}$
-$\quad\times\sqrt{L(L+1)(2L+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}+2\delta_{L,L^{\prime}}\delta_{S,S^{\prime}}(-1)^{L+S^{\prime}+J+1}$
-$\quad\times\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{S(S+1)(2S+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}\bigg{)}^{2}.$ (S5)
+$\Gamma_{\mathrm{M1}}=\frac{\mu_{B}^{2}\omega_{0}^{3}}{3\pi\varepsilon_{0}\hbar c^{5}\!}\,\frac{|\mathcal{M}|^{2}}{2J^{\prime}+1},$ (6)
+where
+$\mathcal{M}$ $=\delta_{S,S^{\prime}}\delta_{L,L^{\prime}}(-1)^{L^{\prime}+S+J^{\prime}+1}\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{L(L+1)(2L+1)}\begin{Bmatrix}L^{\prime}&J^{\prime}&S\\ J&L&1\end{Bmatrix}$ (7)
+$\quad+2\,\delta_{L,L^{\prime}}\delta_{S,S^{\prime}}(-1)^{L+S^{\prime}+J+1}\sqrt{(2J+1)(2J^{\prime}+1)}\sqrt{S(S+1)(2S+1)}\begin{Bmatrix}S^{\prime}&J^{\prime}&L\\ J&S&1\end{Bmatrix}.$
 […]
-Table S1: Parameters of interest for the M1 transition $s^{2}p^{2}\,{{}^{3}}P_{1}\leftrightarrow s^{2}p^{2}\,{{}^{3}}P_{0}$ for all Group IV elements.
+Table 6: Parameters of interest for the M1 transition $s^{2}p^{2}\,{{}^{3}}\!P_{1}\leftrightarrow s^{2}p^{2}\,{{}^{3}}\!P_{0}$ for all Group IV elements.
 $k$ denotes transition wavenumber, $\omega_{0}$ denotes transition energy, $\Gamma_{\rm M1}$ denotes natural linewidth, and $\tau_{g}$ denotes lifetime of the metastable state.
@@
 Lead (Pb) 7819.26 234.42 1.145 0.14
-SM II Further Details on Atomic Tin MOT Simulations
+Appendix B Further Details on Atomic Tin MOT Simulations
 Numerical simulations of all Sn MOT stages used a geometry where the $z$-axis is defined by the direction of the symmetry axis of the quadrupole magnetic field for the MOT.
@@
 These curves are defined by:
-$a(r)=\frac{1}{2v_{\text{max}}}\int_{-v_{\text{max}}}^{v_{\text{max}}}a(r,v)\,dv,$ (S6)
+$a(r)=\frac{1}{2v_{\text{max}}}\int_{-v_{\text{max}}}^{v_{\text{max}}}a(r,v)\,dv,$ (8)
 and:
-$a(v)=\frac{1}{2r_{\text{max}}}\int_{-r_{\text{max}}}^{r_{\text{max}}}a(r,v)\,dr.$ (S7)
+$a(v)=\frac{1}{2r_{\text{max}}}\int_{-r_{\text{max}}}^{r_{\text{max}}}a(r,v)\,dr.$ (9)
 Here, $v_{\text{max}}$ and $r_{\text{max}}$ are the maximum velocity and displacement of atoms in the capture MOT, which we find to be $v_{\text{max}}\approx 8$ m/s and $r_{\text{max}}\approx 3$ mm.
 Plots of these cuts are shown in Fig.
-S1(a) and (b), indicating the trapping and cooling behavior of the capture MOT stage.
+8(a) and (b), indicating the trapping and cooling behavior of the capture MOT stage.
 The sign reversal in the $a(v)$ curve at low velocities is a hallmark of sub-Doppler heating, a signature behavior of Type-II MOTs.
 Fig.
-S1(c) shows how the capture velocity $v_{c}$ is determined, as described in the main text.
+8(c) shows how the capture velocity $v_{c}$ is determined, as described in the main text.
 We find that for velocities up to $v_{c}\sim 28.5$ m/s, atoms entering the MOT are slow enough to be captured.
@@
 Fig.
-S1(d) shows an example phase-space plot illustrative of how the temperature and size of the simulated capture MOT are determined.
+8(d) shows an example phase-space plot illustrative of how the temperature and size of the simulated capture MOT are determined.
 An atom is initialized at the phase-space origin and its trajectory propagated using the calculated acceleration $a(r,v)$.
@@
 We average these values over 10 independent, stochastic trajectories to obtain the reported values for temperature $T$ and cloud size $\sigma$.
-Figure S1: Analysis of the simulated Sn capture MOT.
+Figure 8: Analysis of the simulated Sn capture MOT.
 (a) Acceleration $a$ vs displacement $r$.
@@
 Conveyor belt blue MOT simulations were also analyzed similarly, with emphasis placed on the rms positional compression of the MOT cloud during this stage.
-SM III Further Details on Tin White Light Slowing Simulations
+Appendix C Further Details on Tin White Light Slowing Simulations
 Numerical simulations of white light slowing (WLS) of Sn used a geometry where the atomic beam defines the $z$-axis.
@@
 The time dependence of the laser electromagnetic field is given by:
-$e^{i\omega t+i\beta_{1}\sin(\Omega_{1}t)+i\beta_{2}\sin(\Omega_{2}t)}=e^{i\omega t}\left(J_{0}(\beta_{1})+\sum_{k=1}^{\infty}J_{k}(\beta_{1})e^{ik\Omega_{1}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{1})e^{-ik\Omega_{1}t}\right)\times\\ \left(J_{0}(\beta_{2})+\sum_{k=1}^{\infty}J_{k}(\beta_{2})e^{ik\Omega_{2}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{2})e^{-ik\Omega_{2}t}\right),$ (S8)
+$e^{i\omega t+i\beta_{1}\sin(\Omega_{1}t)+i\beta_{2}\sin(\Omega_{2}t)}=e^{i\omega t}\left(J_{0}(\beta_{1})+\sum_{k=1}^{\infty}J_{k}(\beta_{1})e^{ik\Omega_{1}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{1})e^{-ik\Omega_{1}t}\right)\times\\ \left(J_{0}(\beta_{2})+\sum_{k=1}^{\infty}J_{k}(\beta_{2})e^{ik\Omega_{2}t}+\sum_{k=1}^{\infty}(-1)^{k}J_{k}(\beta_{2})e^{-ik\Omega_{2}t}\right),$ (10)
 where $J_{k}(\beta)$ is the $k$th order Bessel function of the first kind.
@@
 Interpolation across $z$ values from $z=0$ mm to $z=500$ mm enables us to obtain deceleration curves $a_{z}(v_{z},r,z)$, which are shown in Fig.
-S2.
+9.
 The calculated $a_{z}(v_{z},r,z)$ and $a_{\perp}(v_{z},r,z)$ are used to solve the differential equations for the trajectory of Sn atoms as they are slowed from source to MOT region.
-Figure S2: Three-dimensional plot of the longitudinal deceleration $a_{z}(v_{z},z)$ imparted on Sn atoms due to white light slowing.
+Figure 9: Three-dimensional plot of the longitudinal deceleration $a_{z}(v_{z},z)$ imparted on Sn atoms due to white light slowing.
 Profiles for three different transverse displacements from the atomic beam axis $\hat{z}$ are shown: $r=0$ mm (blue), $r=3$ mm (green), and $r=6$ mm (brown).